OpenStudy (anonymous):

What is the solution to the equation? 2- z/3=12 A. z = –42 B. z = –34 C. z = –30 D. z = 42

OpenStudy (anonymous):

Algebra is basically just a balance. You do one thing to one side and you have to do the same to the other. So applying this and the fact that you want to isolate z, you should first take away the 2. To do this you must subtract two from both sides('cause 2 is a positive, so this will cause it to cancel) What do you get when you do this @plzhelp!?!? ?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Good, so -z/3=10 also means z/-3=10. you are dividing z by a -3, so you need to do the opposite and multiply it ON BOTH SIDES. What do you get @plzhelp!?!? ?

OpenStudy (anonymous):

answer is c: z= -30

OpenStudy (anonymous):

@stylestar5678 please don't give answers!

OpenStudy (anonymous):

http://www.wolframalpha.com/widget/widgetPopup.jsp?p=v&id=7014d9722bc42de3f621eaad1da2615e&title=Simplifying%20Algebraic%20Expressions%20Calculator&theme=blue&i0=9x-9y%5E3&podSelect= this website will help you and explain how to break down the problem and solve it step by step

OpenStudy (anonymous):

so you can understand how to get to the final answer

OpenStudy (anonymous):

@GodGirl360 I dont know

OpenStudy (anonymous):

That's fine. Just take (z/-3)-3=(10)-3. It's just like the last step. The -3 on the side z is on will cancel itself, but the equation will still be balanced 'cause you did the same thing on the other side.

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Sorry, no. Think of it this way. the -3 on z's side is being divided so we're going to multiply it and do the same on the other side. This causes the -3 on z's side to completely go away so this is what you get - z=(-3)10. What do you get?

OpenStudy (anonymous):

-30

OpenStudy (anonymous):

sorry that's not suppose to be a negative z.

OpenStudy (anonymous):

Good job! that's the answer. :)

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Welcome! any time.